There is a race to a point 3miles away and back. Person A goes 20 mph, Person B goes 16mph. Where do they meet (in feet).
I figured that it took person A .15 hours to the turning point and the Person B had traveled 2.4 miles in that time.
I don't know how to go the rest of the answer....! Help!

Answer by mananth(16946)

race distance = 3 & back
A = 20 mph
B= 16 mph
..
A takes 3/20 hour to reach other end
..
By then
B has reached 16*3/20 miles
=12/5 miles
..
at this point of time they are 3-(12/5) miles apart facing each other.
=3/5 miles.
..
Now they are running towards each other
their speeds when added up is 36 mph
..
distance = 3/5 miles
speed = 36 mph
..
time they meet = (3/5) /36
3/5 *1/36
=3/180 hours.
..
in 3/180 hours B will run ( d= rt)
3/180 * 20
=1/3 mile
...
1/3 * 5280 = 1760 feet from the returning end.
...
m.ananth@hotmail.ca